OpenStudy (anonymous):

What is the combined perimeter of these two rectangles? A) 6x + 14 B) 12x + 24 C) 12x + 26 D) 24x + 52

OpenStudy (the_beast_eli):

B, I THINK

OpenStudy (texaschic101):

P = 2L + 2W P = 2(2x + 7) + 2(4x + 5) P = 4x + 14 + 8x + 10 P = 12x + 24 P = 2(3x) + 2(3x + 14) P = 6x + 6x + 28 P = 12x + 28 12x + 24 + 12x + 28 = 24x + 52 <=== combined total

OpenStudy (the_beast_eli):

Whoops! Caps locked

OpenStudy (the_beast_eli):

Yeah, I meant that XD

OpenStudy (anonymous):

thx :)

OpenStudy (anonymous):

@texaschic101 im still confused

OpenStudy (anonymous):

@kiamousekia

OpenStudy (texaschic101):

confused about what ?